The KOMA-timedate-bundle^{∗}

Markus Kohm

1995/05/24

Abstract This bundle includes a package scrtime defining some macros to handle compilation-time. It’s a very simple implementation similar to time.sty. I’ve tried to use no additional register.

The second package scrdate defines some macros to handle the name of the day!

Contents

1 Introduction

1.1 scrtime-package

\today

\date {today’s-value} \thistime [hourseparator] \settime {thistime’s-value}

At L^{A}TEX there’s a macro \today to get compilation-date. There’s a macro \date to set the value of \today, too. But to handle compilation-time there is only a macro \time which shows the gone minutes of the day.

With the macro \thistime you get the actual time. The optional argument [hourseparator] is used to separate hours and minutes. Default-value is “:”. \thistime writes values of minutes less than ten with prefix “0”. \thistime* doesn’t.

With the macro \settime you may set the value of \thistime. After this the optional argument of \thistime will be ignored.

1.2 scrdate-package

At L^{A}TEX there are macros to get the actual day or month or year. Using \today you even may get the name of the month, but you will never get the name of the day.

This package defines some macros to handle the name of the day.

∗This file has version number v1.1c, last revised 1995/05/24.

Important: If you use language-packages like german or babel, these must be loaded after scrdate-package. Otherwise the name of the day’ll be wrong!

scrdate knows four languages:

• english

• USenglish

• german

• french

\newdaylanguage {language}{Monday}{Tuesday}{Wednesday}{Thursday}{Friday}{Saturday}{Sunday} You may add additional languages using the configurationfile scrdate.cfg. There you have to define a macro \dayname@language. You may use \newdaylanguage to do so.

\todaysname

shows the name of the actual day (defined by \year, \month and \day or by \dayname).

\dayname {Name}

sets the name of the day to {Name}.